### v2

# Nearly geodesic surfaces are filling

Abstract
Let $M$ be a closed hyperbolic $3$-manifold. A homotopy class $[S]$ of surfaces in $M$ is filling if for any $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components that are contractible in $M$. We prove that there exist $\epsilon_{0},g_{0}>0$ such that every homotopy class of $(1+\epsilon)$-quasi-Fuchsian surfaces with $0<\epsilon\leq\epsilon_{0}$ or totally geodesic surfaces of genus $\geq g_{0}$ in $M$ is filling. As a corollary, except for at most finitely many totally geodesic surfaces, embedded incompressible quasi-Fuchsian surfaces in $M$ have constants bounded below by $1+\epsilon_{0}$. This also gives a gap theorem for embedded minimal surfaces. Each of these surfaces separates any pair of distinct points at $\partial\widetilde{M}$. Crucial tools include the rigidity results of Mozes-Shah, Ratner, and Shah. This work is inspired by a question of Wu and Xue whether random geodesics on random hyperbolic surfaces are filling.
Dedicated to Nathan M. Dunfield with gratitude, in celebration of his 50th birthday, for
his support since 2015.
1 Introduction
Let $M$ be a closed hyperbolic $3$-manifold and $\widetilde{M}$ the universal cover. A surface $S$ in $M$ is $(1+\epsilon)$-quasi-Fuchsian if its limit set $\partial\widetilde{S}\subset\partial\widetilde{M}$ is a $(1+\epsilon)$-quasicircle (but not a $(1+\epsilon^{\prime})$-quasicircle for any $\epsilon^{\prime}<\epsilon$). A homotopy class $[S]$ is filling if for every $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components contractible in $M$. Rubinstein-Sageev [RS99] define $[S]$ to be strongly filling if for any $p\neq q\in\partial\widetilde{M}$, there is an $\widetilde{S}$ covering $S$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S}$. The genus of a totally geodesic surface is the smallest genus of an orientable surface in its commensurability class. In this paper, we prove
Theorem 1.1.
Let $M$ be a closed hyperbolic $3$-manifold. There exist $\epsilon_{0},g_{0}>0$ such that every homotopy class of $(1+\epsilon)$-quasi-Fuchsian surfaces with $0<\epsilon\leq\epsilon_{0}$ or totally geodesic surfaces of genus $\geq g_{0}$ is strongly filling.
Kahn and Marković [KM12] prove that $M$ has infinitely many $(1+\epsilon)$-quasi-Fuchsian surfaces for $\epsilon<\epsilon_{0}$. Moreover, the proof of Theorem 1.1 shows that all but finitely many totally geodesic hypersurfaces of a closed hyperbolic manifold are strongly filling. An immersed surface $S\subset M$ is $\pi_{1}$-injective if the induced homomorphism $\pi_{1}(S)\rightarrow\pi_{1}(M)$ is injective. For the rest of the paper, we set $\epsilon_{0}$ and $g_{0}$ as in Theorem 1.1.
Corollary 1.2.
Let $M$ be a closed hyperbolic $3$-manifold. The set of quasi-Fuchsian $\pi_{1}$-injective embedded surfaces in $M$ is decomposed into
  1. 1.
    at most finitely many totally geodesic surfaces of genus $g_{0}$, and
  2. 2.
    surfaces with quasi-Fuchsian constant $\geq 1+\epsilon_{0}$.
The resolution of the tameness conjecture [Ago04, CG06] implies that a $\pi_{1}$-injective surface in $M$ is either quasi-Fuchsian or virtually fibered (lifts to a fiber of a finite cover that is a surface bundle over $S^{1}$). The supremum principal curvature of a fibered minimal surface $F$ is >1>1 by [Uhl83, HL21] and the Hausdorff dimension of the limit set of $F$ is $2$. Thus 1.2 implies that any sequence of $\pi_{1}$-injective surfaces in $M$ is eventually self-intersecting or uniformly away from being totally geodesic.
Corollary 1.3.
Let $M$ be a closed hyperbolic $3$-manifold. If $S\hookrightarrow M$ is an embedded minimal surface, then $S$ is either a totally geodesic surface of genus $g_{0}$ or a quasi-Fuchsian surface of supreme principal curvature $\geq\frac{\epsilon_{0}}{\epsilon_{0}+2}$.
We say $[S]$ in $M$ separates two distinct points $p,q\in\partial\widetilde{M}$ if $S^{\prime}\in[S]$ has a cover $\widetilde{S^{\prime}}\subset\widetilde{M}$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S^{\prime}}$. Kahn-Marković [KM12] constructs a collection of nearly geodesic surfaces which is dense in the Grassmannian $2$-plane bundle $\mathcal{G}_{2}M$ and Bergeron-Wise [BW12, Corollary 4.2] points out that these nearly geodesic surfaces satisfy a boundary criterion.
Theorem 1.4 (KM12).
Let $M$ be a closed hyperbolic $3$-manifold. Every pair of distinct points $p,q\in\partial\widetilde{M}$ is separated by some quasi-Fuchsian $[S]$ in $M$.
This theorem, along with other important results of Agol, Bergeron-Wise, Haglund-Wise, Sageev, and Wise, has important consequences in geometric group theory and virtual properties of $3$-manifolds, see Section 3.6, [Ago13, HW08, Sag95, Wis21], and the references therein. The boundary criterion of Bergeron-Wise provides a motivation to study strongly filling surfaces .
If $[S]$ is strongly filling, then for any $^{\prime}\in[S], SS^{\prime}$ intersects every homotopically nontrivial closed curve in $M$, which implies that $[S]$ is filling. By 2.7, a least area representative of a filling homotopy class actually cuts the ambient manifold into topological balls. In Section 3.6 we point out that a strongly filling $[S]$ gives a cubulation with a single orbit of hyperplanes (whose stabilizer is essential). Non commensurable strongly filling surfaces give rise to cubulations that are distinct under the isometric actions by $\pi_{1}(M)$ and thus their existence answers a question of Fioravanti-Hagen on [FH21, p. 879].
Corollary 1.5.
Let $M$ be a closed hyperbolic $3$-manifold or a closed arithmetic hyperbolic manifold with a totally geodesic hypersurface. The space of essential and hyperplane-essential cubulations with a single orbit of hyperplanes whose hyperplane-stabilizers are hypersurface subgroups is infinite up to subdivision and natural isometric actions by $\pi_{1}(M)$.
Figure 1: The filling nature of nearly geodesic surfaces in a closed hyperbolic $3$-manifold (each dot represents a $(1+\epsilon)$-quasi-Fuchsian surface with $\epsilon\leq\epsilon_{0}$).
Before talking about the proof of Theorem 1.1, we summarize it in Figure 1, which illustrates the idea that a certain neighborhood of totally geodesic surfaces of genus $\geq g_{0}$ in the space of all $\pi_{1}$-injective surfaces consists of filling surfaces. In various arguments, it is helpful for the author to view transverse intersections and filling as an “open” property of $\mathcal{G}_{2}M$ and of the space of $\pi_{1}$-injective surfaces in $M$. This figure is inspired by [KMS23, Figure 1] and justified in Section 2.1.
Two $\pi_{1}$-injective surfaces $S_{1},S_{2}\subset M$ are commensurable if there exists a $g\in\pi_{1}(M)$ such that $g\pi_{1}(S_{1})g^{-1}\cap\pi_{1}(S_{2})$ is of finite index in both $\pi_{1}(S_{1})$ and $\pi_{1}(S_{2})$. A commensurability class of surfaces is filling if each of its homotopy classes is filling. We always assume surfaces are orientable and $\pi_{1}$-injective unless stated otherwise. A surface is Fuchsian or totally geodesic if any geodesic on $S$ with respect to its induced Riemannian metric is also a geodesic on $M$. Commensurable Fuchsian surfaces correspond to the same $PSL_{2}(\mathbb{R})$-orbit. A sequence of $(1+\epsilon)$-quasi-Fuchsian minimal surfaces is asymptotically Fuchsian if $\epsilon\rightarrow 0$. The proof of Theorem 1.1 uses Theorem 1.6 and 3.15.
Theorem 1.6.
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{j}$ is a sequence of asymptotically Fuchsian surfaces in distinct commensurability classes, then the Grassmann $2$-plane bundle $\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}M$ in the Hausdorff metric.
Remark 1.7.
Independently, Fernando Al Assal and Ben Lowe [AL25] obtain Theorem 1.6 for finite-volume hyperbolic manifolds with different perspectives and applications. Al Assal [AA23] shows that the measure limits of $\mathcal{G}_{2}S_{i}$ achieve every convex combination of $SL_{2}(\mathbb{R})$-invariant ergodic measures.
Vladimir Marković suggested the “one versus uncountably many planes duality” for the limiting behavior of asymptotically Fuchsian surfaces to the author, which led to Theorem 1.6. 3.15 then establishes filling by the following argument. Suppose there exist nontrivial $\gamma_{i}]\in\pi_{1}(M-S_{i})$. Then a subsequence of $(x_{i},v_{i})$ in the unit tangent bundle $T^{1}\gamma_{i}$ of the geodesic $\gamma_{i}$ converges to $(x,v)$ tangent to a geodesic $\gamma$. There exist $(p_{j},\Pi_{j})\in\mathcal{G}_{2}S_{j}\rightarrow(x,\Pi)\perp(x,v)$. Since $\epsilon_{j}\rightarrow 0$, large $S_{j}$ are nearly orthogonal to $\gamma_{j}$, and then $\partial\tilde{S}_{j}$ separates $\partial\widetilde{\gamma_{j}}$ which means $\pi_{1}(M-S_{j})$ cannot contain $\gamma_{j}$. We need to use [Sep16] which controls the norms of the second fundamental forms of $\tilde{S}_{j}$ by $\epsilon_{j}$ and [Smi07] to prove $\tilde{S}_{j}$ converge to a totally geodesic plane $\widetilde{P}$.
A geodesic plane in $M$ is the image of a totally geodesic plane $\widetilde{P}\subset\mathbb{H}^{3}\rightarrow M$. The proof of Theorem 1.6 uses crucially [Rat91, Sha91a] which imply a geodesic plane in $M$ is either dense or a closed Fuchsian surface, and uses [MS95] which implies that a sequence of Fuchsian surfaces in distinct commensurability classes equidistribute (become uniformly dense in $\mathcal{G}_{2}M$). Since $S_{i}$ are minimal, their Grassmann bundles are connected closed sets, and their Hausdorff limit is a connected closed subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. Any finite union of the Grassmann bundles over Fuchsian surfaces is a disconnected closed set. Thus, $\mathcal{G}_{2}S$, which is closed and $PSL_{2}(\mathbb{R})$-invariant, must contain a dense plane in $\mathcal{G}_{2}M$.
Theorem 1.1 is inspired by Theorem 1.8. Given a Fuchsian $S$, let $\mathbf{g}_{c}(S)$ be the smallest genus of an orientable surface in $M$ commensurable with $S$.
Theorem 1.8.
Let $M$ be a closed hyperbolic $3$-manifold. Given a sequence of Fuchsian surfaces $S_{i}\hookrightarrow M$ with $\mathbf{g}_{c}(S_{i})\rightarrow\infty$, the maximal diameter of components of $M-S_{i}$ tends to 0 and the number of components tends to infinity.
The proof of Theorem 1.8 actually shows that a sequence of totally geodesic hypersurfaces cuts the ambient closed hyperbolic $n$-manifold ($n\geq 3$) into components whose diameters tend to 0. Since $M$ is aspherical, a component of MSM-S of diameter $2\textnormal{inj}(M)$ is necessarily contractible. By [BFMS21] (for $n\geq 3$) and [MM22] (for $n=3$), if $M$ contains infinitely totally geodesic hypersurfaces, then $M$ is arithmetic.
An outline of the proof of Theorem 1.8 is the following. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Suppose there exist components $C_{i}\subset M-S_{i}$ with $\textnormal{diam}(C_{i})\geq 2R$. Since $\pi^{-1}(C_{i})\subset\mathbb{H}^{3}$ consists of convex components, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length =2R=2R. Up to a subsequence, $\gamma_{i}\rightarrow$ a geodesic segment $\gamma$ of length $2R$. 3.7 proves that the subset of $\mathcal{G}_{2}M$ whose points are tangent to the planes transversally intersecting the central half of $\gamma$ with uniformly bounded angles contains an open set of $\mathcal{G}_{2}M$ of uniformly positive measure (depending only on $R$). Thus eventually $S_{i}$ must intersect the central half of $\gamma$ with uniformly bounded angles. This shows that $S_{i}$ must transversally intersect $\gamma_{i}$ for large $i$, a contradiction. If the maximal diameter of every component of $M-S_{i}$ is less than $2\textnormal{inj}(M)$, then $S_{i}$ is filling. [LM23] points out the rate of equidistribution and hence $g_{0}$ depends only on the rate of mixing, $\textnormal{Vol}(M)$, and the injectivity radius. 3.18 proves that a filling Fuchsian surface necessarily has some transverse intersections with all geodesics in $M$, which we are uncertain to hold for filling least area non Fuchsian surfaces.
One way to visualize Theorem 1.8 and its transition to Theorem 1.1 is the following. Let $D$ be a convex polyhedron fundamental domain for $M$ in $\mathbb{H}^{3}$. The generators of $\pi_{1}\rightarrow\Gamma\leq PSL_{2}(\mathbb{C})$ are given by finitely many side-pairings of $\partial D$, i.e., $\{g\in\Gamma:g\overline{D}\cap\overline{D}\subset\partial\overline{D}\textnormal{ and}\neq\emptyset\}$. The preimages of large filling Fuchsian surface give rise to some geodesic planes that cut $D$ and $\partial D$ into components so that any geodesic that passes through $D$ must intersect one of the planes transversally.
In addition to being an $SL_{2}(\mathbb{R})$-orbit in $\mathbb{H}^{3}$, a geodesic plane can be defined as a submanifold with principal curvatures 0. A slight perturbation to these geodesic planes gives rise to minimal planes with small principal curvatures and limit sets with small quasi-Fuchsian constants. Since the polyhedron fundamental domain is compact, finitely many evenly distributed almost-geodesic planes remain candidates for filling. In other words, although the proof of Theorem 1.8 relies on the rigidity of $SL_{2}(\mathbb{R})$-orbits, the topological filling nature persists for nearby surfaces which are not homogeneous (i.e., orbits of Lie subgroups of $SL_{2}(\mathbb{C})$).
A stronger notion of separation than filling is ubiquitous. A collection of immersed surfaces in a hyperbolic 33–manifold $M$ is ubiquitous if, for any pair of disjoint and nonasymptotic planes $\Pi,\Pi^{\prime}\subset\mathbb{H}^{3}$, some surface in the collection has an embedded preimage $\tilde{S}\subset\mathbb{H}^{3}$ that separates $\Pi$ from $\Pi^{\prime}$. Kahn and Marković [KM12] prove that in a closed hyperbolic $3$-manifold, the set of nearly geodesic surfaces they construct is ubiquitous. Cooper and Futer [CF19] prove that in a cusped hyperbolic $3$-manifold, there exists a set of $K$-quasi-Fuchsian surfaces which is ubiquitous for some $K>1$, while Kahn and Wright [KW21] prove that for any $K>1$, there exists a ubiquitous set of $K$-quasi-Fuchsian surfaces. Al Assal [AA23] proves the existence of asymptotically Fuchsian surfaces, based on [LM15, KW21]. Theorem 1.6 implies the following.
Corollary 1.9.
Let $M$ be a closed hyperbolic $3$-manifold. A sequence of asymptotically Fuchsian surfaces in distinct commensurability classes is ubiquitous.
The proofs of 1.9 and Theorem 1.6 rely on the rigidity of unipotent flow, i.e., theorems of Ratner [Rat91] and Shah [Sha91a], which is very different from [KM12, CF19, KW21] (but 1.9 and Theorem 1.6 do not give the existence of nearly geodesic surfaces). The analogous statement to 1.9 for cusped hyperbolic $3$-manifolds is a corollary of the main result of [AL25].
Although the property of being filling sorts out a subclass of submanifolds of a hyperbolic $3$-manifold, it also reflects certain topological complexity of the ambient manifold. Recall from Thurston [Thu97, Definition 3.8.1] that a geometric $3$-manifold is a Riemannian $3$-manifold supporting one of the eight model geometries. Nathan Dunfield first suggested the existence of $\pi_{1}$-injective filling surfaces (with infinite fundamental groups) in a non hyperbolic geometric $3$-manifold, i.e., the Hantzsche-Wendt manifold.
Theorem 1.10.
If $M$ is a geometric $3$-manifold and admits a $\pi_{1}$-injective filling homotopy class of surfaces, then $M$ is hyperbolic, spherical, Euclidean, or $\mathbb{H}^{2}\times\mathbb{R}$.
In the proof we explicitly construct some filling surfaces; see Figure 5 for the Euclidean manifolds and Figure 6 for the $\mathbb{H}^{2}\times\mathbb{R}$ ones. We also introduce an elementary inequality connecting the volume of a hyperbolic $3$-manifold with the area of its filling surfaces. By 2.7, a least area surface in a filling $[S]$ cuts $M$ into balls and thus naturally $\textnormal{Area}(S)$ and $\textnormal{Vol}(M)$ satisfy the isoperimetric inequality for $M$, which was pointed out by Wu and Xue [WX22]. The hyperbolic space can be characterized as a space satisfying linear isoperimetric inequalities. Let $\textnormal{rank}(M)$ be the smallest number of generators of $\pi_{1}(M)$.
Theorem 1.11.
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling homotopy class of $\pi_{1}$-injective surfaces. Then any representative $S$ satisfies
(1.1) $\textnormal{Area}(S)>\textnormal{Vol}(M),$
and there is an absolute constant $C$ such that
(1.2) $\textnormal{rank}(S)>C\cdot\textnormal{rank}(M).$
Current estimates give $C$ an order of $10^{-9}$. The qualitative linearity between the rank of $S$ and $M$ in (1.2) contrasts other geometries: 4.4 gives a sequence of $3$-manifolds $M_{j}$ with $\mathbb{H}^{2}\times\mathbb{R}$ geometry such that $\textnormal{rank}(M_{j})\rightarrow\infty$ but all admit a filling torus. The proof of (1.1) is a natural consequence of the linear isoperimetric inequality for hyperbolic $3$-space. The proof of (1.2) relies on (1.1) and [BGLS10] which bound the rank from below by the volume for symmetric space. Cooper [Coo99] bounds the volume by the presentation length which inspires (1.2).
We summarize Table 1 for convenience.
Properties 22-dim 33-dim
Minimizing measures of a hypersurface minimizes its intersection number with closed geodesics. False in general, only true for geodesic submanifolds (3.2 and 3.22).
Filling geodesic hypersurfaces have some transverse intersections with every geodesic. for hyperbolic $n$-manifolds with $n\geq 3$ (3.18).
Random geodesic hypersurfaces are filling (a conjecture originated from Wu and Xue [WX22]). Open for random Weil-Petersson model (partial results for expander surfaces by [DS23]). in every closed hyperbolic manifold of dim $\geq 3$, for all but finitely many commensurability classes (Theorem 1.1).
There exist infinitely many embedded geodesic submanifolds. No, not even asymptotically geodesic, 1.2.
The support of measure convergence is distinct from that of Hausdorff convergence. for many sequences of geodesics by Thurston ([Thu97, Mar16]). for asymptotically Fuchsian surfaces ([AL25] and Theorem 1.6, [AL25]) unless eventually commensurable.
Table 1: The beautiful analogies and contrasts between the submanifold theory of hyperbolic surfaces and that of hyperbolic $3$-manifolds.
The motivations for Theorem 1.1 are twofold: one from the statistics of the topological properties of closed geodesics on a random hyperbolic surface, in particular, the work of Wu and Xue [WX22], and the other from the rigidity of unipotent flows.
1.1 The motivations from hyperbolic geometry and random geodesics
The polar coordinate for the $2$-dimensional hyperbolic metric is
(1.3) $g=\,\mathrm{d}r^{2}+\sinh^{2}r\,\mathrm{d}\theta^{2}.$
A disk of radius $r$ has an area of $2\pi(\cosh r-1)$ and a perimeter of $2\pi\sinh r$. Let $B$ be a compact contractible domain $\subset\mathbb{H}^{2}$ with piecewise smooth boundary $\partial B$. The linear isoperimetric inequality (e.g., [Bus10, p. 211]) is
$\textnormal{Area}(B)<\ell(\partial B).$
Since a genus-$g$ hyperbolic surface $S_{g}$ has an area of $4\pi(g-1)$, its fundamental domain $D\subset\mathbb{H}^{2}$ has a boundary length of at least $4\pi(g-1)$, and thus a filling geodesic must have a length of at least $2\pi(g-1)$. On $S_{g}$, there are only finitely many closed geodesics of length less than $2\pi(g-1)$. Wu and Xue in [WX22, Question on p. 5] ask the following.
Question 1.12.
As $g\rightarrow\infty$, on a generic closed hyperbolic surface $S$ in the Weil-Petersson model, are most closed geodesics of length significantly greater than $g$ filling?
Dozier and Sapir in [DS23, Section 5.1] construct a sufficiently fine triangulation $\mathcal{D}$ of $S$ and flow boxes centered at the edges of $\mathcal{D}$ so that most sufficiently long geodesics pass through all these boxes and hence are filling. Their ideas apply to hypersurfaces in higher-dimensional manifolds and detect filling. However, for Theorems 1.8 and 1.1, we have adapted an argument different from [DS23] to prove large Fuchsian surfaces are filling by taking advantage of the rigidity of $SL_{2}(\mathbb{R})$-invariant orbits explained below. The crucial difference is that, given any prescribed subset of $\mathcal{G}_{2}M$, all large Fuchsian surfaces must get close and nearly tangent to.
1.2 The motivations from the rigidity of unipotent flows
There is a huge variety of invariant sets of geodesic flows on the unit tangent bundle $T^{1}S$ of a hyperbolic surface, i.e., dense [KH95, 17.5, Corollary 18.3.5], closed geodesics, geodesic laminations [Mar16, 8.3], and sets with Hausdorff dimension greater than $1$ but less than $3$ [LS18, Theorem 1.3]. The flexibility of such invariant sets is also reflected by the existence of arbitrarily long closed geodesics supported on a proper subsurface of $S$, which are never filling. Although Fuchsian surfaces have principal curvature 0 and geometrically generalize geodesics, it is their connection with dynamic systems that gives rise to the filling topological nature in Theorem 1.8 and Theorem 1.1. Calegari-Marques-Neves [CMN22] initiated the studies of the distribution of minimal surfaces in negatively curved manifolds using homogeneous dynamics and rigidity results.
Let $G$ be a connected Lie group and $U$ a unipotent one-parameter subgroup of $G$. In contrast with geodesic flows, there is a strong rigidity for the orbits of horocycle flow ($U$-orbits): Hedlund’s theorem [Hed36] proves the horocycle flow on a closed hyperbolic surface is minimal, i.e., every horocycle on a closed hyperbolic surface is dense. Shah [Sha91a] classifies the orbit closure of SO(n1,1)SO(n-1,1) for actions on quotients of $SO(n,1)$ for $n\geq 3$; Etienne Ghys pointed out the implications in geometry: the projection of a codimension-one geodesic hyperplane $P\subset\mathbb{H}^{n}$ to a closed hyperbolic $n$-manifold $M$ is either dense or a closed totally geodesic hypersurface. For $n\geq 3$, the stark contrast between SO(n1,1)SO(n-1,1) and geodesic $SO(1,1)$ relies crucially on the behavior of the orbits of nontrivial unipotent $1$-parameter subgroups contained in SO(n1,1)SO(n-1,1), as can be seen in [Sha91a, Section 4-6]. Margulis [Mar89] shows how to effectively study orbit closures in homogeneous spaces of Lie groups by analyzing minimal closed invariant sets of the action of unipotent subgroups. Some other important works include, for example, Dani-Smillie, Dani, Dani-Margulis, which we refer the readers to the references in e.g., [Mar89, Sha91a, Sha91b, Rat91].
Ratner [Rat91] classifies finite ergodic $U$-invariant measures and proves every unipotent subgroup of a connected Lie group $G$ is strictly measure rigid, which establishes Raghunathan’s measure conjecture. Shah proves a theorem [Sha91b, 1.1] which describes geometrically a subset of points whose $U$-orbits are uniformly distributed with respect to a $G$-invariant measure; in particular, combining with the measure classification [Rat91], he in [Sha91b, 1.3] verifies a stronger form of Raghunathan’s orbit closure conjecture for cocompact quotients of $G$. [MS95] implies that a sequence of Fuchsian surfaces $S_{i}$ in distinct commensurability classes becomes equidistributed, i.e., the probability measure induced by the inclusion of $S_{i}$ on the Grassmann bundle $\mathcal{G}_{2}M$ converges to the uniform Liouville measure on $\mathcal{G}_{2}M$, building on [Rat91]. There is a rapid development of effective Ratner’s theorem. Readers are referred to [LM23, LMW22, Yan22] and the references therein.
Lowe [Low21] studies the minimal foliations of Grassmann bundles of negatively curved manifolds. He also proves various quantitative density and measure convergence results for asymptotically Fuchsian surfaces in non-homogeneous metrics, e.g., [Low21, Theorem 1.9]. Jiang [Jia25] generalizes the minimal surface entropy results of [CMN22] to higher-dimensional and noncompact finite-volume hyperbolic manifolds. Thurston in [Thu97, Remark after Corollary 8.8.6] points out that unlike closed hyperbolic $3$-manifolds, a noncompact finite-volume hyperbolic manifold might have infinitely many homotopy classes of $\pi_{1}$-injective surfaces of bounded genus. Jiang [Jia25] proves that the number of homotopy classes of nearly geodesic surfaces of bounded genus remains finite. Lowe [BLM24], following ideas of Labourie [Lab21], Lowe-Neves [LN21], proves that the weak- limit of the probability measures induced by a sequence of asymptotically Fuchsian surfaces is $PSL_{2}(\mathbb{R})$-invariant.
An outline of the paper is the following. In Section 2, we recall some background on quasi-Fuchsian and minimal surfaces, and the metric and measure on Grassmann bundles. In Section 3, we prove Theorems 1.8 and 1.1, 1.2, 1.3, and 1.5. In Section 4, we classify geometric $3$-manifolds containing a filling $\pi_{1}$-injective surface, construct explicit filling surfaces in Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ geometry, and prove Theorems 1.10 and 1.11.
Acknowledgment
The author would like to thank Wenyuan Yang for suggesting a connection between Theorem 1.8 and the work of Kahn-Marković. He thanks Vladimir Marković for insightful discussions on the asymptotic behaviors of the planes which help formulate Theorem 1.6 and part of Theorem 1.1. He also thanks Nathan M. Dunfield, Mark Hagen, and Zhenghao Rao for many helpful discussions. He thanks his postdoc mentor Yunhui Wu for many helpful discussions in the last few years on hyperbolic geometry and minimal surfaces. Part of the work is completed when the author visits the Institut Fourier. The author would like to thank the institute for its hospitality and thank Andrea Seppi and Zeno Huang for their interest in this work and for helpful discussions.
2 Preliminaries
2.1 Quasi-Fuchsian surfaces and minimal surfaces
We call $S\subset M$ primitive if $\pi_{1}(S)$ is not a proper subgroup of the fundamental group of another orientable subsurface. Since $M$ is hyperbolic, a Fuchsian surface $S$ of genus $g$ has intrinsic sectional curvature 1-1. The conjugacy classes of surface subgroups $\pi_{1}(S)$ of $\pi_{1}(M)$ are in one-to-one correspondence with the homotopy classes of $\pi_{1}$-injective immersions $f:S\rightarrow M$. We define a local variation of $f(S)$ to be a smooth map
$F:S\times(-\eta,\eta)\rightarrow M\quad(\eta>0)$
with $F(x,0)=x$ for all $x\in S.$ For sufficiently small $|t|$, $\Psi_{t}(\cdot)\coloneqq F(\cdot,t)$ defines a smooth immersion of $S$ into $M$. A surface $S$ is minimal if
$\frac{d}{dt}\textnormal{Area}(\Psi_{t}(S))|_{t=0}=0$
for every local variation $\Psi$. By [SU82, SY79], a $\pi_{1}$-injective immersion $f:S\rightarrow M$ is homotopic to a smoothly immersed minimal surface. Henceforth, we assume all such immersions $f:S\rightarrow M$ are minimal unless stated otherwise. A $\pi_{1}$-injective surface $S$ is least area if $S$ achieves the smallest area in its homotopy class. For further background on minimal surfaces, see [CM11] and [Jos17, 5.4].
An immersed surface $f:S\hookrightarrow M$ is $(1+\epsilon)$-quasi-Fuchsian if the limit set of a universal cover $\tilde{S}\subset\mathbb{H}^{3}$ is a $(1+\epsilon)$-quasicircle (but not a $(1+\epsilon^{\prime})$-quasicircle for any $\epsilon^{\prime}<\epsilon$). Commensurable surfaces share the same quasi-Fuchsian constants.
Seppi proves the following theorem.
Theorem 2.1 (Sep16, Theorem A).
There exist universal positive constants $\epsilon_{1}$ and $C$ such that for every $(1+\epsilon)$-quasi-Fuchsian minimal surface with $\epsilon\leq\epsilon_{1}$, its supremum principal curvature $\lambda_{0}\coloneqq\|\lambda\|_{L^{\infty}}$ satisfies
(2.1) $\lambda_{0}\leq C\epsilon.$
Conversely, Z. Huang and B. Wang prove
Theorem 2.2 (HW13, (3.16)).
If a $(1+\epsilon)$-quasi-Fuchsian minimal surface $S$ satisfies $\lambda_{0}<1$, then
$1+\epsilon<\frac{1+\lambda_{0}}{1-\lambda_{0}}.$
Both $\epsilon=0$ and $\lambda_{0}=0$ correspond to totally geodesic surfaces. In a hyperbolic manifold, by the Gauss-Bonnet theorem and the Gauss equation, the area $T$ of a minimal surface $S$ is $\leq 4\pi(g-1)$, with equality if and only if $S$ is totally geodesic. Moreover, combining with Theorem 2.1, we can derive that the area of a $(1+\epsilon_{0})$-quasi-Fuchsian surface is $\geq(\frac{1}{1+C^{2}\epsilon_{0}^{2}})4\pi(g-1)$. This justifies Figure 1.
A minimal surface with $\lambda_{0}<1$ is almost-Fuchsian. Uhlenbeck [Uhl83] pioneered the studies of such surfaces, using the metric and the second fundamental form of an almost-Fuchsian surface in a quasi-Fuchsian $3$-manifold $M$ to parameterize the metric tensor of $M$. An unpublished result of Thurston states that almost-Fuchsian surfaces are $\pi_{1}$-injective, see [Lei06, Theorem 5.1].
Lemma 2.3 ([Uhl83] and [Sep16]).
Let $M$ be a closed hyperbolic $3$-manifold. There exists an $\epsilon_{2}>0$ (depending only on $M$) such that for $\epsilon<\epsilon_{2}$, every $(1+\epsilon)-quasi-Fuchsian$ surface is homotopic to a unique minimal surface.
Throughout this work, we assume $\epsilon_{0}$ (e.g., in Theorem 1.1) satisfies $\epsilon_{0}\leq\min\{\epsilon_{1},\epsilon_{2}\}$. For recent advances on almost-Fuchsian manifolds and surfaces, see [HLS23] and the references therein.
When analyzing filling properties and geodesic intersections, there is no loss of generality in considering orientable surfaces of minimal genus within each class. On a hyperbolic surface $S$, a primitive closed geodesic $\gamma$ and its cover $\gamma^{n}$ span identical subsets of $S$. Similarly, a least area primitive surface $S$ and other surfaces commensurable to it span identical subsets of $M$, and their preimages in $\mathbb{H}^{3}$ are the same by [FHS83, Theorem 5.3]. A hypersurface of a Riemannian manifold is a closed codimension-$1$ submanifold with the induced Riemannian metric.
2.2 A natural metric on the Grassmann bundles
Following [Pat99], we construct a natural metric on the Grassmann $2$-plane bundle for the convenience of the readers. A point in $\mathcal{G}_{2}M$ is denoted by $(x,\Pi)$ where $x\in M$ and $\Pi\subset T_{x}M$ is a $2$-plane. The unit tangent bundle $T^{1}M$ is a $2$-fold covering of $\mathcal{G}_{2}M$.
For this subsection only, let $\pi:\mathcal{G}_{2}M\rightarrow M$ be the canonical projection, i.e., if $\theta=(x,\Pi_{1})\in\mathcal{G}_{2}M$, then $\pi(\theta)=x$. There exists a canonical subbundle of $T\mathcal{G}_{2}M$ called the vertical subbundle: its fiber at $\theta$ is given by the tangent vectors of curves $\sigma:(-\eta,\eta)\rightarrow\mathcal{G}_{2}M$ of the form $\sigma(t)=(x,\Pi_{1}+t\Pi_{2})$, where $\Pi_{2}\in(\mathcal{G}_{2})_{x}M$. Equivalently,
$V(\theta)=\ker(d_{\theta}\pi).$
We now use the Riemannian metric on $M$ to construct a metric on $\mathcal{G}_{2}M$ and a complementary horizontal subbundle. There is a horizontal lift
$L_{\theta}:(\mathcal{G}_{2})_{x}M\rightarrow T_{\theta}\mathcal{G}_{2}M$
using the parallel transport. Given $\Pi\in(\mathcal{G}_{2})_{x}M$ and $\alpha:(-\eta,\eta)\rightarrow M$ such that $\alpha^{\prime}(0)\perp\Pi$ in $T_{x}M$, let $Z(t)$ be the parallel transport of $\Pi$ along $\alpha$. Let $\sigma:(-\eta,\eta)\rightarrow\mathcal{G}_{2}M$ be the curve
$\sigma(t)=(\alpha(t),Z(t)).$
Then
$L_{\theta}(\Pi)\coloneqq\sigma^{\prime}(0)\in T_{\theta}\mathcal{G}_{2}M.$
The horizontal subbundle is defined as $\textnormal{Im}(L_{\theta})$. By [Pat99, Lemma 1.15], we see that as vector spaces
$T_{\theta}\mathcal{G}_{2}M=H(\theta)\oplus V(\theta).$
Equip each fiber with $\frac{1}{2\pi}$ times the standard Riemannian metric on $\mathbb{R}P^{2}$, so that the volume measure of each fiber $(\mathcal{G}_{2})_{x}$ is $1$.
Definition 2.4 (A natural invariant metric on the Grassmann bundle $\mathcal{G}_{2}M$).
Using the decomposition $T_{\theta}\mathcal{G}_{2}M=H(\theta)\oplus V(\theta)$, we define the Sasaki metric on $\mathcal{G}_{2}M$ by declaring $H(\theta)\oplus V(\theta)$ to be orthogonal. Denote this metric by $\langle\cdot\,,\cdot\rangle_{\mathcal{G}_{2}M}$.
The orthogonal splitting endows $\mathcal{G}_{2}M$ with a natural metric $\langle\cdot\,,\cdot\rangle_{\mathcal{G}_{2}M}$ satisfying the following properties:
  1. 1.
    the map $\pi:\mathcal{G}_{2}M\rightarrow M$ is a Riemannian submersion with totally geodesic fibers,
  2. 2.
    the projection of a geodesic in $\mathcal{G}_{2}M$ to $M$ is a geodesic, in particular,
    $\textbf{{d}}_{\mathbb{H}^{3}}(x_{1},x_{2})\leq\textbf{{d}}_{\mathcal{G}_{2}\mathbb{H}^{3}}((x_{1},\Pi_{1}),(x_{2},\Pi_{2})),$
    and
  3. 3.
    the lift of this metric to $\mathcal{G}_{2}(\mathbb{H}^{3})$ is invariant under $\textnormal{Isom}(\mathbb{H}^{3})$, and agrees with hyperbolic metric on totally geodesic planes in $\mathcal{G}_{2}(\mathbb{H}^{3})$.
The Sasaki metric induces a probability measure on $\mathcal{G}_{2}M$ called the Liouville measure, invariant under the geodesic flow.
Let $f:S\rightarrow M$ be an $\pi_{1}$-injective quasi-Fuchsian surface. There is a hyperbolic metric on $S$ with respect to which $f$ is a conformal minimal immersion. Here $f$ also induces a natural map to $\mathcal{G}_{2}M$, which we still denote by $f$.
Definition 2.5 (Probability measure induced by minimal surfaces).
For a measurable subset $O\subset\mathcal{G}_{2}M$, and $f:S\rightarrow\mathcal{G}_{2}M$, we have
$\mu_{S}(O)=\frac{\textnormal{Area}(f^{-1}(O\cap\mathcal{G}_{2}S))}{\textnormal{Area}(S)}.$
2.3 Fundamental domains and filling
We recall some results about fundamental domains of a hyperbolic manifold $M$, which help us visualize how a filling surface interacts with geodesics and $M$.
By [Rat19, Theorem 6.6.13], a discrete subgroup of $\textnormal{Isom}(\mathbb{H}^{n})$ admits a locally finite Dirichlet fundamental domain $D$. By [Rat19, Exercise 6.6.2], the discrete group $\Gamma=\pi_{1}(M)$ is generated by side-pairings:
$\{g\in\Gamma:\bar{D}\cap g\bar{D}\subset\partial\overline{D}\textnormal{ and}\neq\emptyset\}.$
By the first paragraph of [Rat19, 6.7], there exists a convex $D$. Furthermore, [Rat19, Theorem 6.7.1] proves that $\bar{D}$ is a convex polyhedron. Finally, [Rat19, Theorem 6.7.4(2), Theorem 6.7.5] prove that the fundamental domain $D$ is exact: for any side $F$ of $D$, there exists a unique $g\in\Gamma$ such that
$F=\bar{D}\cap g\bar{D}.$
This culminates in the following lemma.
Lemma 2.6.
A closed hyperbolic $n$-manifold admits a fundamental domain $D\subset\mathbb{H}^{n}$ which is a convex polyhedron, whose sides are partitioned into pairs $\{\{F_{1},F_{1}^{\prime}\},\cdots,\{F_{m},F_{m}^{\prime}\}\}$, such that for each pair of sides $F_{i},F_{i}^{\prime}$, there exists a unique $g_{i}\in\Gamma$ which maps $F_{i}$ isometrically onto $F_{i}^{\prime}$. Moreover, the group $\Gamma$ is generated by $\{g_{1},\cdots,g_{m}\}$.
We call a surface $S\hookrightarrow M$ filling if MSM-S consists of contractible components. We always assume surfaces in hyperbolic $3$-manifolds are quasi-Fuchsian. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Then $\pi(\partial D)$ is a filling closed surface which is homotopically trivial. This motivates us to define a filling homotopy class of surfaces as a property that should hold for every representative [^1] We thank Chao Wang for clarifications regarding filling surfaces versus filling homotopy classes.. A filling homotopy class $[S]$ intersects every homotopically nontrivial loop of $M$ which means $\pi_{1}(M-S)\rightarrow\pi_{1}(M)$ is trivial. We call a disconnected space contractible if each of its components is contractible. We record the following useful lemma from Rubinstein-Sageev.
Lemma 2.7 (RS99).
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling homotopy class of surfaces. Then a least area representative $S$ is filling.
Thus similar to the boundary of a fundamental domain, a least area surface in a filling homotopy class has an area lower bound by the linear isoperimetric inequality (see Theorem 1.11).
3 Strongly filling surfaces in hyperbolic $3$-manifolds
We start with some lemmas on Fuchsian and filling surfaces, whose proofs also help us visualize the proof of Theorem 1.1.
3.1 Some lemmas on intersections and filling
We first define the intersection number between $\gamma]\in\pi_{1}(M)$ and a homotopy class $[S]$ of immersed surfaces. By applying homotopies, we may assume an immersed closed curve $\gamma:S^{1}\rightarrow M$ and an immersed surface $f:S\rightarrow M$ intersect at finitely many points $x_{1},\cdots,x_{n}$, such that for each $i\in\{1,\cdots,n\}$, there exist $j\in\{1,\cdots,i(\gamma)\}$, $k\in\{1,\cdots,i(S)\},t_{ij}\in S^{1}$, and $y_{ik}\in S$ with
$\gamma(t_{ij})=x_{i}=f(y_{ik})\textnormal{ and }\gamma^{\prime}(t_{ij})\oplus df_{y_{ik}}(T_{y_{ik}}S)=T_{x_{i}}M.$
The intersection number of $\gamma$ and $S$ is defined by counting with multiplicity:
$|\gamma\cap S|=\sum_{i=1}^{n}i(\gamma)\cdot i(S)$
Definition 3.1.
The intersection number between a homotopy class $\gamma]\in\pi_{1}(M)$ and a homotopy class $[S]$ of surfaces is
$\gamma]\cdot[S]\coloneqq\min\{|\gamma^{\prime}\cap S^{\prime}|:\gamma^{\prime}\in[\gamma],S^{\prime}\in[S]\}.$
If a Fuchsian surface is filling, it must self-intersect sufficiently to intersect all of its own geodesics transversally. A Fuchsian surface minimizes the intersection number with a closed geodesic in its homotopy class.
Lemma 3.2.
Let $\gamma]\in\pi_{1}(M)$ and $[S]$ be a homotopy class of a Fuchsian surface. Then the intersection number $\gamma]\cdot[S]$ is realized by the number of transverse intersections of their geodesic representatives $\gamma$ and $S$
$\gamma]\cdot[S]=|\gamma\cap S|.$
Proof.
Lift $M$ to a fundamental domain $D\subset\mathbb{H}^{3}$. The preimages of $\gamma$ and $S$ in $D$ are finitely many geodesic arcs and geodesic polygons. If an arc which is a subset of a geodesic line $\tilde{\gamma}$ has a transverse intersection with a totally geodesic plane $P$, then the circle $\partial P$ separates the two endpoints of $\tilde{\gamma}$. By the Morse lemma, any homotopy of $\gamma$ induces a homotopy of $\tilde{\gamma}$, fixing its limit points. Similarly, any homotopy of $S$ induces a homotopy of $P$, fixing the circle at infinity. Thus no homotopy decreases the intersection number between $\beta$ and $P$. Applying this argument to all transverse intersections between the preimages of $\gamma$ and the preimages of $S$ through $D$, we conclude that the minimal number is realized by taking the geodesic representatives of both $\gamma$ and $S$. ∎
We call the intersection between a Fuchsian surface and a closed geodesic in a hyperbolic manifold essential. The strong interplay between topology and geometry is a recurring theme in hyperbolic geometry. If $M=S^{2}\times S^{1}$ and $\gamma\subset S^{2}$ a geodesic great circle, then $\gamma\times S^{1}\subset M$ is a compressible totally geodesic torus. Thus 3.2 fails in general.
As a corollary of 3.2, we have
Lemma 3.3.
If a Fuchsian surface $S$ intersects a geodesic $\gamma$ transversally, then for any $S^{\prime}$ homotopic to $S$, $\pi_{1}(M-S^{\prime})$ does not contain $\gamma]$. Moreover, if a Fuchsian surface $S$ is filling, its homotopy class and commensurability class are strongly filling.
Thus if a Fuchsian $S$ transversally intersects every closed geodesic of $M$, $[S]$ is filling. This local geometric criterion for a homotopy class of geodesic surfaces to be filling also generalizes to certain non hyperbolic manifolds, e.g., the proof of 4.3.
Remark 3.4.
3.2 and 3.3 do not generalize to least area non-Fuchsian surfaces in hyperbolic manifolds: 3.22 proves every such surface has redundant intersections with some closed geodesics.
Lemma 3.5.
Let $S\subset M$ be a filling surface and $M^{\prime}$ is a cover of $M$. Then $S$ lifts to a (possibly disconnected or noncompact) filling surface $S^{\prime}\subset M^{\prime}$.
Proof.
Denote by $\pi:M^{\prime}\rightarrow M$ the covering projection. If $S^{\prime}=\pi^{-1}(S)$ is not filling, then there exists a non-contractible component $B$ of $M^{\prime}-S^{\prime}$. Then $\pi(B)\subset M-S$ is non-contractible, which is a contradiction. ∎
Let $\mathcal{G}_{2}M$ be the Grassmann $2$-plane bundle over $M$. We denote by
  1. 1.
    $\pi:\mathbb{H}^{3}\rightarrow M$ the universal covering map, and
  2. 2.
    $\textnormal{Pr}:\mathcal{G}_{2}M\rightarrow M$ the natural projection.
Define the set of planes in $\mathcal{G}_{2}M$ by
(3.1) $\mathcal{P}(M)\coloneqq\mathcal{G}_{2}M/\sim$
[…]
Definition 3.6.
Two commensurable closed Fuchsian surfaces are tangent to the same discrete plane. The genus of a discrete plane $P$ is the smallest genus of the orientable Fuchsian surfaces tangent to $P$.
This is the genus $g_{c}(S)$ as in Theorem 1.8. Since there are only finitely many Fuchsian surfaces of genus up to $g$, and no distinct commensurability classes of Fuchsian surfaces are tangent, the set of planes up to genus $g$ is a disconnected closed set in $\mathcal{G}_{2}M$ (when this set has more than one planes). A subset $B$ of $\mathcal{G}_{2}M$ is filling if $\textnormal{Pr}(A)\subset M$ is filling. For $x\in M$ and $\Pi_{x}\in\mathcal{G}_{2}M$ tangent to a plane $P$, we call $\Pi_{x}$ intersects a geodesic $\gamma$ transversally in $M$ if $\gamma$ and $P$ intersect at some point transversally. Thus for a filling Fuchsian surface $S$, any $x\in S$, and any geodesic $\gamma\subset S$, $\Pi_{x}$ intersects $\gamma$ transversally by 3.18.
3.2 Large Fuchsian surfaces are strongly filling
One way to see that as the genus of discrete planes grows, they must cut $M$ into more and more components is as follows. Using 2.6, $M$ admits a fundamental domain $D\subset\mathbb{H}^{3}$, which is a compact convex polyhedron of diameter $d$. The intersection of $D$ with any geodesic plane in $\mathbb{H}^{3}$ has an area upper bound $2\pi(\cosh\frac{d}{2}-1)$ using the polar coordinates (1.3). Thus as the area of $S$ grows, the lift of $S$ in $D$ consists of more and more geodesic polygons in $D$. [MS95] implies that these polygons are uniformly distributed and thus must cut $M$ into components of decreasing diameter.
For $R>0$, define two positive functions
$h_{1}(R)=\cot^{-1}\left(\frac{\sinh\frac{R}{4}\cosh\frac{R}{2}}{\sinh\frac{R}{2}}\right)-\cot^{-1}\left(2\cosh^{2}\frac{R}{4}\right)$
and
$h_{2}(R)=\frac{\pi(1-\cos\theta_{0})}{4}(R\cosh R-R).$
Lemma 3.7.
Let $s$ be a geodesic segment of length $2R\leq 2\textnormal{inj}(M)$. Then the open subset of $\mathcal{G}_{2}M$ whose points intersect the central half of $s$ transversally with an angle $h_{1}(R)$ has Liouville measure $\geq h_{2}(R)$.
Proof.
Let $s=s(z)$ be a unit-speed parameterization where $z\in I^{\prime}=[-R,R]$ and $s(I^{\prime})$ is embedded. Let $I=[\frac{-R}{2},\frac{R}{2}]$ parameterize the central half of $s$. Extend $z$ to the Fermi coordinates $(r,\theta,z)$ centered at $s$, where $r,\theta$ form the $2$-dimensional polar coordinate for the hyperbolic disk orthogonal to $s$ at $s(z)$. The hyperbolic metric is given by $\textnormal{d}r^{2}+\sinh^{2}r\,\textnormal{d}\theta^{2}+\cosh^{2}r\,\textnormal{d}z^{2}$. Let $\phi_{1},\phi_{2}$ be the spherical coordinates for the unit tangent space at $(r,\theta,z)$, where $\phi_{1}\in[0,\pi]$ measures the angle relative to $z$-axis and $\phi_{2}\in[0,2\pi]$ corresponds to $\theta$. Hence $(r,\theta,z,\phi_{1},\phi_{2})$ form local coordinates for $(x,v)$ in the unit tangent bundle $T^{1}M$ which double covers $\mathcal{G}_{2}M$.
We now describe an open subset $O(s)\subset\mathcal{G}_{2}M$ whose points are tangent to planes transversally intersecting $s(I)$ with angle $h_{1}(R)$. Using the ball model, we place the segment $s$ at a vertical $\mathbb{H}^{2}$ so that its perpendicular bisector passes through the origin. Then $s(I)=\{(0,0,z)|z\in[-R/2,R/2]\}$. The angle $\theta=0$ (resp. $\theta=\pi$) corresponds to the right (resp. left) half-space of the complement of the geodesic line containing $s(I)$ in $\mathbb{H}^{2}$. The coordinates $r\in[0,R/2]$ and $\theta\in[0,2\pi]$ parameterize a solid annulus $A$ centered at $s$. Let $\theta_{0}$ be the angle formed by the segment $\{(r,\pi,R/4)|r\in[0,R/2]\}$ and the segment $(R/2,\pi,R/4),(0,0,R/2)]$. By applying [Bus10, (i), (iii), (vi)] to the right triangle with side-length $a=R/4,b=R/2,\alpha=\theta_{0}$, and $\beta$, we see
$\theta_{0}=\cot^{-1}\left(2\cosh^{2}\frac{R}{4}\right)\textnormal{ and }\beta=\cot^{-1}\left(\frac{\sinh\frac{R}{4}\cosh\frac{R}{2}}{\sinh\frac{R}{2}}\right).$
At $x=(r,\theta,z)\in A$, let $v$ be the unique unit vector orthogonal to a plane which is orthogonal to $s$. Let $O(\theta_{0})\subset T^{1}_{x}M$ be vectors forming an angle with $v$ less than or equal to $\theta_{0}$. Then $O(\theta_{0})$ is a disk on the unit sphere with measure $\frac{1-\cos\theta_{0}}{2}$. The plane through $x$ orthogonal to $(\phi_{1},\phi_{2})=(\theta_{0},\pi)$ intersects $s(I)$ at an angle $\beta$ and the plane orthogonal to $(\theta_{0},\pi/2)$ intersects $s(I)$ at an angle $\pi/2-\theta_{0}$. We deduce that every plane orthogonal to a vector in $O(\theta_{0})$ intersects $s(I)$ at an angle
$\cot^{-1}\left(\frac{\sinh\frac{R}{4}\cosh\frac{R}{2}}{\sinh\frac{R}{2}}\right)-\cot^{-1}\left(2\cosh^{2}\frac{R}{4}\right).$
Thus $O(s)$ has measure $\geq$
$\int_{0}^{R/2}\int_{0}^{2\pi}\int_{-R/4}^{R/4}(1-\cos\theta_{0})\sinh r\cosh r\,\textnormal{d}z\,\textnormal{d}\theta\,\textnormal{d}r=\frac{\pi(1-\cos\theta_{0})}{4}(R\cosh R-R).$
Using the metric on the Grassmann bundle (2.4), any $(x,\Pi)\in O(s)$ is $\frac{\theta_{0}}{\sqrt{2\pi}}$ close to (one which is tangent to a plane) orthogonal to $s$.
The proof of Theorem 1.8.
Proof.
Recall $\mathbf{g}_{c}(S)$ denotes the minimal genus of an orientable surface commensurable with $S$. We first prove that the maximal diameter of components of $M-S_{i}$ tends to 0 as $g_{i}\coloneqq\mathbf{g}_{c}(S_{i})\rightarrow\infty$. Suppose, for contradiction, that there exists a subsequence $S_{i}$ with a component $C_{i}\subset M-S_{i}$ such that $\textnormal{diam}(C_{i})\geq 2R>0$ (we may assume $R<\textnormal{inj}(M)$).
Since $S_{i}$ is totally geodesic, each component of $\mathbb{H}^{3}-\pi^{-1}(S_{i})$ is the intersection of finitely many geodesic half-spaces and thus must be convex. In particular, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length $2R$. Since $M$ is compact, up to taking a subsequence, $\gamma_{i}$ converges to a geodesic segment $\gamma$ of length $2R$. Using 3.7, the subset of Grassmann bundles tangent to planes transversally intersecting the central half of $\gamma$ with an angle $\geq h_{1}(R)$ contains an open set $O$ of measure $\mu_{\mathcal{L}}(O)\geq h_{2}(R)>0$.
By [MS95], the surfaces $S_{i}$ become equidistributed, i.e., for $\mu_{i}\coloneqq\mu_{S_{i}}$
$\mu_{i}\rightharpoonup\mu_{\mathcal{L}}$
in the weak-$^{*}$ topology, where $\mu_{\mathcal{L}}$ is the Liouville measure supported on $\mathcal{G}_{2}M$. Thus
$\liminf\mu_{i}(O)\geq\mu_{\mathcal{L}}(O)\geq h_{2}(R)>0.$
This implies that for sufficiently large $i$, $S_{i}$ intersects the central half of $\gamma$ with an angle $\geq h_{1}(R)$. Combining this with the Hausdorff convergence $\gamma_{i}\rightarrow\gamma$, we conclude that $S_{i}$ must intersect $\gamma_{i}$ transversally. Contradiction.
Since the maximal diameter of components of $M-S_{i}$ tends to 0, their maximal volume also tends to 0. Thus the number of components necessarily tends to infinity. ∎
If a closed hyperbolic $n$-manifold $M^{n}$ contains a sequence of totally geodesic hypersurfaces $S_{i}$, since $S_{i}$ equidistribute by [MS95], the proof of Theorem 1.8 similarly shows that $S_{i}$ cuts $M^{n}$ into components whose diameters tend to 0. Each component $O$ of $M-S_{i}$ lifts to a convex subset of $\mathbb{H}^{n}$. If $\textnormal{diam}(O)<2\textnormal{inj}(M)$, then $O$ is contractible in $M$. This finishes the proof of the statement from Theorem 1.1: there exists a constant $A_{0}(M)$ such that every totally geodesic hypersurface whose Hausdorff measure $\geq A_{0}(M)$ is filling.
Given a closed hyperbolic $3$-manifold with exact generators $\{g_{1},\cdots,g_{m}\}$ of $\pi_{1}(M)\rightarrow\Gamma$ (in the sense of 2.6) for a fundamental domain $D$ with sides $F_{i},F_{i}^{\prime}$ (for $i=1,\cdots,m$), one can prove there exist $2m$ open subsets $O_{i},O_{i}^{\prime}$ of $\mathcal{G}_{2}M$ so that a Fuchsian $S$ whose $\mathcal{G}_{2}S$ passes through all $O_{i}$ and $O_{i}^{\prime}$ are filling, using 3.7. Each $O_{i}$ (resp. $O_{i}^{\prime}$) corresponds in $M$ to nearly orthogonal to $g_{i}$ and close to $F_{i}$ (resp. $F_{i}^{\prime}$) in the Grassmann bundle metric $d_{\mathcal{G}_{2}M}$.
We obtain the following corollary on the set $\mathcal{P}(M)$ of planes. Denote $\mathcal{P}_{\textnormal{fill}}(M)=\{P\in\mathcal{P}(M):M-P\textnormal{ consists of contractible components}\}.$
Corollary 3.8.
The subset $\mathcal{P}_{\textnormal{fill}}(M)$ of filling planes is the dense open set $\mathcal{P}(M)-\sqcup_{i=1}^{m}P_{i}$ where $P_{i}$ corresponds to some non-filling surface.
Proof.
By Ratner’s [Rat91] and Shah’s [Sha91a] theorem, a plane corresponds to either a dense subset of $\mathcal{G}_{2}M$ or a closed Fuchsian surface. The projection of a dense subset of $\mathcal{G}_{2}M$ must be filling in $M$, and there are at most finitely many commensurability classes of non-filling Fuchsian surfaces by Theorem 1.1. ∎
Remark 3.9.
The above proof and hence Theorem 1.8 are valid for higher-dimensional hyperbolic manifolds and their totally geodesic hypersurfaces.
A closed hyperbolic $3$-manifold lies in exactly one of the following categories:
  1. 1.
    arithmetic with infinitely many commensurability classes of Fuchsian surfaces (by [Rei91]),
  2. 2.
    arithmetic with no Fuchsian surfaces (by [Rei91]), or
  3. 3.
    non-arithmetic with at most finitely many commensurability classes of Fuchsian surfaces (by [BFMS21]).
Theorem 1.8 trivially holds for the second and third categories. [MR03, Theorem 9.5.4] shows how to construct infinitely many arithmetic hyperbolic $3$-manifolds containing Fuchsian surfaces. We first proved Theorem 1.8 and that large Fuchsian surfaces are filling. Vladimir Marković further suggested that most nearly geodesic surfaces Jeremy Kahn and he constructed in [KM12] should be filling, which led to part of the formulation of Theorem 1.1.
3.3 Nearly geodesic surfaces are strongly filling
[AA23] and the proofs of Theorem 1.6 and Theorem 1.1 will reveal the subtle difference between measure convergence and Hausdorff convergence. It is illuminating to examine hyperbolic surfaces first. Let $S$ be a genus-$2$ hyperbolic surface and $\beta$ a simple close geodesic. By $D_{\beta}$ we mean applying the Dehn twist with respect to $\beta$. Let $\mathcal{PML}$ denote the space of projective measured laminations on $S$ and $C(S)$ denote the collection of closed sets in $S$ equipped with Hausdorff metric. From the basic theory on geodesic laminations (e.g., [Mar16, 8.3]), we have as $n\rightarrow\infty$,
$\frac{1}{2n}D_{\beta_{1}}^{n}D_{\beta_{2}}^{n}\gamma$ $\rightarrow\frac{1}{2}\beta_{1}\sqcup\frac{1}{2}\beta_{2}\textnormal{ in }\mathcal{PML}\textnormal{ with the weak-}^{*}\textnormal{ topology,}$
$D_{\beta_{1}}^{n}D_{\beta_{2}}^{n}\gamma$ $\rightarrow\lambda_{1}\cup\beta_{1}\cup\lambda_{2}\cup\beta_{2}\textnormal{ in }C(S)\textnormal{ with the Hausdorff metric topology,}$
where $\lambda_{i}$ are simple complete geodesic spiraling around $\beta_{1},\beta_{2}$. See Figure 2.
(a) Applying Dehn twists to $\gamma$ with respect to $\beta_{1},\beta_{2}$.
(b) The Hausdorff limit carries a connected line field.
Figure 2: The weak- limit is supported on two components. The Hausdorff limit connects the two components by two geodesics which are neither closed nor dense and forms a connected subset of the line bundle over $S$.
In contrast, every geodesic plane in a closed hyperbolic $3$-manifold $M$ is either dense or a closed Fuchsian surface by Ratner’s and Shah’s theorem [Rat91, Sha91a]. In Al Assal’s thesis [AA23], he proves that the set of weak- limits of the probability area measures induced on the Grassmann bundle by asymptotically Fuchsian minimal surfaces achieves every convex combination of ergodic $SL_{2}(\mathbb{R})$-invariant measures. When $M$ has no Fuchsian surfaces, the limiting measure must be the Liouville measure which implies that $\mathcal{G}_{2}S_{i}$ equidistribute. When $M$ has a Fuchsian surface, Theorem 1.6 suggests that there is a dense subset in the Hausdorff limit with measure 0. In particular, to connect distinct (or identical) components in the Hausdorff limit, we must connect by a dense subset, unless the asymptotically Fuchsian surfaces belong to the same commensurability class (in which case the geodesic planes in $\mathbb{H}^{3}$ or $\mathcal{G}_{2}M$ which cover them coincide).
Lemma 3.10.
A $\pi_{1}$-injective surface $S$ in a $3$-manifold is homotopic to a smoothly immersed minimal surface whose Grassmann bundle is a connected closed subset of $\mathcal{G}_{2}M$.
Proof.
The existence of a smooth minimal surface representative follows from Schoen-Yau [SY79, Lemma 3.2 and Theorem 2.1] and Sacks-Uhlenbeck [SU82]. Osserman [Oss70] and Gulliver [Gul73] prove such minimal surfaces have no branched points. The smoothness of the tangent bundle of $S$ implies that $\mathcal{G}_{2}S$ is a connected closed subset of $\mathcal{G}_{2}M$. The uniqueness of the minimal surface representative in 2.3 allows us to identify a homotopy class $[S]$ of $(1+\epsilon)$-quasi-Fuchsian surfaces with a unique subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. ∎
[AA23] proves that for a sequence of asymptotically Fuchsian surfaces $S_{i}$, both the pleated maps $f^{p}_{i}$ and minimal maps $f^{m}_{i}$ homotopic to $S_{i}\hookrightarrow M$ have the same weak- measure limit along any convergent subsequence. The universal cover of $f^{p}_{i}$ may develop wrinkles and each $f^{p}_{i}$ lifts to a disconnected subset of $\mathcal{G}_{2}M$.
We now prove the following theorem which implies Theorem 1.1.
Theorem 3.11.
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{i}$ is a sequence of asymptotically Fuchsian surfaces in distinct commensurability classes, then all but finitely many $S_{i}]$ are strongly filling.
Proof.
We proceed by contradiction. Suppose $S_{i}]$ is a sequence of $(1+\epsilon_{i})$-quasi-Fuchsian surfaces that are not strongly filling, where $\epsilon_{i}\rightarrow 0$. For each $S_{i}]$, we take the minimal surface representative $S_{i}$. After taking a subsequence and renaming it if necessary, we have
Lemma 3.12.
The Grassmann bundles $\mathcal{G}_{2}S_{i}$ converge to a connected closed $PSL_{2}(\mathbb{R})$-invariant subset of $\mathcal{G}_{2}M$.
Proof.
Let $C(\mathcal{G}_{2}M)$ denote the set of closed subsets in $\mathcal{G}_{2}M$, equipped with Hausdorff metric. Since $M$ is closed, $\mathcal{G}_{2}M$ is closed, and hence $C(\mathcal{G}_{2}M)$ is a compact metric space. Thus $\mathcal{G}_{2}S_{i}$ has a convergent subsequence, which we still denote by $\mathcal{G}_{2}S_{i}$. By 3.10, each $\mathcal{G}_{2}S_{i}$ is a connected closed subset of $\mathcal{G}_{2}M$, and thus
$\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}S\subset\mathcal{G}_{2}M$
[…]
Here $S$ is either one or uncountably many geodesic planes in $\mathbb{H}^{3}$.
Lemma 3.13.
If the limit $\mathcal{G}_{2}S$ is not a single geodesic plane which is discrete, it is dense and $\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}M$.
Proof.
Enumerate the collection of discrete geodesic planes as $\{P_{i}\}$ which are embedded and mutually disjoint in $\mathcal{G}_{2}M$. Denote
$\mathcal{P}_{j}$ $\coloneqq\{P_{i}:i\leq j\},$
$\mathcal{P}_{L}$ $\coloneqq\{\textnormal{geodesic planes which are dense in }\mathcal{G}_{2}M\}.$
For any finite $j$, the union $\cup_{\mathcal{P}_{j}}P_{i}$ is a disconnected closed set in $\mathcal{G}_{2}M$. Hence $\mathcal{G}_{2}S\cap\cup_{\mathcal{P}_{j}}P_{i}$ is a strict subset of $\mathcal{G}_{2}S$ for any $j$, unless $\mathcal{G}_{2}S\subset P_{i}$ for some $i$, contradiction. Thus for any genus $g$, there exist planes $P_{k}$ of genus >g>g such that $P_{k}\subset\mathcal{G}_{2}S-\cup_{\mathcal{P}_{j}}P_{i}$ by 3.12. Since $P_{k}$ equidistribute as $k\rightarrow\infty$ by [MS95], $P_{k}\subset\mathcal{G}_{2}S$ pointed converge to a dense plane $P\in\mathcal{P}_{L}$. Since $\mathcal{G}_{2}S$ is closed and $PSL_{2}(\mathbb{R})$-invariant, $P\subset\mathcal{G}_{2}S$. ∎
Now suppose that $\mathcal{G}_{2}S$ is supported on a single discrete geodesic plane $P$. Then for any small $\eta>0$, $\mathcal{G}_{2}S_{i}\subset N_{\eta}(\mathcal{G}_{2}S)$ for all sufficiently large $i$.
Lemma 3.14.
Let $S\subset M$ be totally geodesic and $\eta$ smaller than the normal injectivity radius of $\mathcal{G}_{2}S$ in $\mathcal{G}_{2}M$. If $S^{\prime}$ is smoothly immersed $\pi_{1}$-injective and $\mathcal{G}_{2}S^{\prime}\subset N_{\eta}(\mathcal{G}_{2}S)$, then $S^{\prime}$ and $S$ have homotopic images.
Proof.
By a homotopy, $S^{\prime}$ is minimal. Since $\mathcal{G}_{2}S$ and $N_{\eta}(\mathcal{G}_{2}S)$ are embedded, the preimages of $N_{\eta}(\mathcal{G}_{2}S)$ in $\mathcal{G}_{2}\mathbb{H}^{3}$ consist of pairwise disjoint embedded neighborhoods $\{N_{\eta}(P)|\textnormal{Pr}(P)\subset\mathbb{H}^{3}\textnormal{ cover }S\}$ (where $P$ is a totally geodesic plane). By 3.10, $\mathcal{G}_{2}S^{\prime}$ is connected, and hence lifts to $\mathcal{G}_{2}\widetilde{S}^{\prime}\subset N_{\eta}(P)$. The projection $\textnormal{Pr}:\mathcal{G}_{2}\mathbb{H}^{3}\rightarrow\mathbb{H}^{3}$ maps $\mathcal{G}_{2}\widetilde{S}^{\prime}$ to $\widetilde{S}^{\prime}\subset N_{\eta}(\textnormal{Pr}(P))$. The only complete minimal disk contained in $N_{\eta}(\textnormal{Pr}(P))$ is the totally geodesic plane $\textnormal{Pr}(P)$ itself. Thus $\widetilde{S}^{\prime}=\textnormal{Pr}(P)$ and $S^{\prime}$ is commensurable with $S$. ∎
Lemma 3.15.
If a sequence of Grassmann bundles $\mathcal{G}_{2}S_{i}$ over minimal surfaces converges to a dense subset of $\mathcal{G}_{2}M$, then all but finitely many $S_{i}$ are filling.
Proof.
[…]
We just proved there is a subsequence $S_{j}$ of the initial sequence $S_{i}$ which is strongly filling. If there exist infinitely many non-filling $S_{i}$, then we apply the above argument again to this collection and conclude it must contain a subsequence consisting of filling surfaces, which is a contradiction. ∎
3.13 establishes Theorem 1.6. Lowe in [Low21, Theorem 4.2] proves that the asymptotically Fuchsian surfaces are dense in $\mathcal{G}_{2}M$ assuming they have a limiting circle which is not the limit set of a closed Fuchsian surface. Fernando Al Assal and Ben Lowe [AL25] independently obtain Theorem 1.6 for finite-volume hyperbolic manifolds (including the noncompact case) and obtain various applications, including rigidity results for negatively curved manifolds [AL25, Theorem 1.7] and a gap theorem [AL25, Theorem 1.6] for the quasi-Fuchsian constants of nearly geodesic surfaces in geometrically finite hyperbolic manifolds of infinite volume.
Remark 3.16.
Al Assal in [AA23] constructs asymptotically Fuchsian surfaces which equidistribute or whose limit has a nontrivial scarring component. Building upon Liu-Markovic [LM15], in [AA23, Section 4, 5, and 7], in order to construct connected asymptotically Fuchsian surfaces with various limiting properties, Al Assal uses all $(R,\epsilon)$-good pants which are known to equidistribute (different limiting measures correspond to different weights on the set of good pants). Thus 3.13 and [AL25] imply as $\epsilon$ gets smaller, the pants required to construct a connected $(1+\epsilon)$-quasi-Fuchsian surface become dense.
Kahn, Marković, and Smilga [KMS23] study the average limiting measures supported by the nearly geodesics surfaces. They discover that the distribution of random surfaces with bounded genus is drastically different from that of random surfaces with bounded area.
Since the proof of Theorem 1.6 relies on the fact that asymptotically Fuchsian surfaces eventually become dense in $\mathcal{G}_{2}M$, this naturally implies that they form a ubiquitous set.
The proof of 1.9.
Proof.
The idea is essentially in the proof of 3.15. Using Theorem 1.6, the Grassmann bundles $\mathcal{G}_{2}S_{i}$ converge to $\mathcal{G}_{2}M$. Let $C_{1},C_{2}$ be two disjoint and non tangent circles at $\partial\widetilde{M}$ which bound two geodesic planes $P_{1},P_{2}\subset\mathbb{H}^{3}$, respectively. We want to show there exists some $S_{i}$ which has a preimage $\tilde{S}_{i}$ whose limit set separates $C_{1}$ and $C_{2}$.
[…]
Then the second paragraph of the proof of 3.15 implies that for sufficiently large $i$, $\partial\tilde{S}_{i}$ separates $C_{1}$ and $C_{2}$ and $\tilde{S}_{i}$ separate $P_{1}$ and $P_{2}$. ∎
3.4 The dichotomy of embedded surfaces in a hyperbolic $3$-manifold
In this subsection, we prove the dichotomy for embedded surfaces in 1.2. The following lemma is due to Feustel [Feu70, Theorem 1], Jaco [Jac71, Theorem 5], and Scott [Sco72, Theorem 1.4]. We present a concise proof with more modern terminology for the convenience of the readers.
Lemma 3.17.
Let $M$ be a compact $3$-manifold. If $S$ is a closed $\pi_{1}$-injective surface which nontrivially covers another orientable $\pi_{1}$-injective surface $S^{\prime}$ embedded in $M$, then $S$ is not embedded.
Proof.
The incompressibility of $S$ and $S^{\prime}$ induces an injection of the fundamental group:
$\pi(S)\rightarrow\pi_{1}(S^{\prime})\rightarrow\pi_{1}(M).$
Suppose $S$ is embedded in $M$. Then, by lifting to the covering $M^{\prime}$ of $M$ corresponding to $\pi_{1}(S^{\prime})$, we also lift $S$ homeomorphically to an embedded incompressible surface $\tilde{S}$ in $M^{\prime}$. By the classification of $I$-bundles over surfaces [Hem76, 10.6 Theorem], we have $M^{\prime}\cong S^{\prime}\times I$. By the classification of incompressible embedded surfaces in an $I$-bundle [Mar16, Proposition 9.3.18], $\tilde{S}$ is isotopic to the horizontal surface $S^{\prime}\times 0$. This contradicts the assumption that $\tilde{S}$ is a nontrivial covering of $S^{\prime}$. ∎
Proof of 1.2.
Proof.
Let $S$ be an embedded surface in $M$. By 3.17, $S$ is a primitive orientable surface. By the Lefschetz duality (e.g., [Mar16, Corollary 9.1.5]), we have $\dim H^{1}(M-S;\mathbb{Q})>0$ which implies that $S$ is never filling. By Theorem 1.1, there exists an $\epsilon_{0}>0$ such that every $(1+\epsilon)$-quasi-Fuchsian surface with $0<\epsilon\leq\epsilon_{0}$ is filling. Thus an embedded non Fuchsian surface satisfies $\epsilon>\epsilon_{0}$.
Now let $S\hookrightarrow M$ be an embedded minimal surface of supremum curvature $\lambda_{0}\leq\frac{\epsilon_{0}}{2+\epsilon_{0}}<1$. We want to show $S$ is totally geodesic of genus $g_{0}$. By Uhlenbeck [Uhl83], having supremum curvature <1<1 implies that $S$ is almost-Fuchsian, say $(1+\epsilon)$-quasi-Fuchsian. By Theorem 2.2,
$1+\epsilon<\frac{1+\lambda_{0}}{1-\lambda_{0}}\leq 1+\epsilon_{0}.$
If $\epsilon\neq 0$, then Theorem 1.1 forces $S$ to be filling and hence non embedded. Thus $\epsilon=0$ and $S$ is Fuchsian. Suppose the genus $\mathbf{g}(S)\geq g_{0}$. If $S$ is an orientable surface attaining the smallest genus in its commensurability class, then Theorem 1.8 implies $S$ is filling. If $S$ is commensurable with a Fuchsian surface of genus $g_{0}$, since commensurable Fuchsian surfaces span the same subset of $M$ (i.e., geometrically identical or same $PSL_{2}(\mathbb{R})$-orbit), the immersion $S\hookrightarrow M$ maps some points in $S$ to points in $M$ with multiplicity more than $1$. This violates the assumption that $S$ is embedded. Thus $S$ must have genus $g_{0}$. ∎
[Miy94, Theorem 4.2] bounds the number of embedded Fuchsian surfaces in $M$ by the volume of $M$ up to a uniform constant. Let $\mathcal{E}_{0}$ be the set of embedded Fuchsian surfaces. There are many non-Haken manifolds $M$ in which case $\mathcal{S}^{E}=\emptyset$. A uniform quasi-Fuchsian constant lower bound suggests most embedded surfaces are some definite amount away from totally geodesic. This can also be measured geometrically using a uniform lower bound on the supremum principle curvature $\lambda_{0}$, i.e.,
$\inf\limits_{S\in\mathcal{E}_{\geq\epsilon_{0}}}\lambda_{0}(S)\geq\delta_{1}$
when we take $S$ to be minimal in its homotopy class. The principal curvature lower bound uses [HW13, Theorem 1.2]. A fibered surface is uniformly away from being totally geodesic by considering either the supremum principal curvature [Uhl83] or the Hausdorff dimension of its limit set.
In general, it is rare that a hyperbolic $3$-manifold contains a Fuchsian surface. Basilio, Lee, and Malionek [BLM24] develop an algorithm that detects whether an ideally triangulated $3$-manifold contains Fuchsian surfaces: out of 150,000 manifolds, they find $9$.
3.5 A filling geodesic surface is strongly filling
Recall that a homotopy class $[S]$ of surfaces in a hyperbolic manifold $M$ separates $p,q\in\partial\widetilde{M}$ if any $S^{\prime}\in[S]$ has a cover $\widetilde{S}$ such that $\partial\widetilde{S}$ separate $p,q\in\partial\widetilde{M}$. Contractible surfaces in $M$ can be homotoped to be filling. However, [RS99, Lemma 2.1] proved that if a Fuchsian surface $S$ is filling, its homotopy class $[S]$ is strongly filling. We present a slightly simpler proof.
Lemma 3.18.
Let $M$ be a closed hyperbolic $n$-manifold. If a totally geodesic hypersurface $S\subset M$ is filling, it has some transverse intersections with every geodesic in $M$. In particular, a filling totally geodesic hypersurface is strongly filling.
Proof.
Let $S$ be a filling totally geodesic hypersurface. By a plane of $\mathbb{H}^{n}$, we mean a totally geodesic hyperplane. Since $S$ is immersed, each pair of planes in $\mathbb{H}^{n}$ which cover $S$ has either empty or transverse intersections. Suppose
$M-S=\cup_{i=1}^{n}B_{i},\textnormal{ where }B_{i}\textnormal{ is open and contractible},$
[…]
Thus $\gamma\cap\partial B_{i}\neq\emptyset$ for some $i$. If $\gamma$ transversally intersects the interior of some polyhedron $F_{j}$ transversally, by lifting, we find a plane which separates $\partial\tilde{\gamma}$.
Each $F_{j}$ is compact and contractible. If $\gamma\subset F_{j}$, since $\gamma$ is complete, $\gamma$ must intersect an (n2)(n-2)-dimensional polyhedron $\alpha\subset\partial F_{j}$. However, $\partial F_{j}$ is a subset of the self-intersection locus of $S$, which means there is another face $F_{k}$ so that $\alpha\subset F_{j}\cap F_{k}$. Thus $\gamma$ intersects $F_{k}$ transversally. By lifting, we find a plane separating $\partial\tilde{\gamma}$. ∎
The idea of the proof also shows that a totally geodesic hypersurface $S\hookrightarrow M$ is filling if and only if every component of $\widetilde{M}-\pi^{-1}(S)$ is bounded.
Remark 3.19.
The argument above relies on the analysis of the intersection of a geodesic $\gamma$ and a geodesic face $\subset$ a totally geodesic plane $P$ and the dichotomy that either $\tilde{\gamma}\subset\widetilde{P}$ or $\tilde{\gamma}\pitchfork\widetilde{P}$. This gives a geometric criterion for a homotopy class of totally geodesic surfaces to be filling in certain aspherical manifolds. See Section 4.1 for filling homotopy classes of surfaces in Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ 33-manifolds.
3.6 Cubulations and one-orbit of hyperplanes
Throughout this subsection, let $X$ be a CAT(0) cube complex. We define an equivalence relation on the edges of $X$ generated by identifying opposite edges of some square in $X$. Given an equivalence class $[e]$ of edges, the hyperplane $w$ dual to $[e]$ is the collection of midcubes which intersect edges in $[e]$. The complement XwX-w consists of two halfspaces. Let $G$ be a finitely generated group with a Cayley graph $\Delta$. A subgroup $H\leq G$ is codimension-$1$ if it has a finite neighborhood $N_{r}(H)$ such that $\Delta-N_{r}(H)$ contains at least two $H$-orbits of deep components, which do not lie in any $N_{s}(H)$. A CAT(0) cube complex $X$ is essential, if for each hyperplane $w$, each of the associated halfspaces contains points in $X$ arbitrarily far from $w$. If a group $G$ acts by automorphism on $X$, the action is essential if for any point $x$ in the zero-skeleton of $X$, and each hyperplane $w$, each of the associated halfspaces contains points in $G\cdot x$ arbitrarily far from $w$. The cube complex $X$ is hyperplane-essential if each hyperplane $w$, regarded itself as a CAT(0) cube complex, is essential. The action of $G$ on $X$ is hyperplane-essential if each hyperplane $w$ has the property that the stabilizer of $w$ acts essentially on $w$.
For more backgrounds and basic definitions of cube complex and its connection to $3$-manifolds, we refer the readers to [AFW15, FH21, BW12] and the references therein. Sageev [Sag95] introduced a method to construct a proper action on a CAT(0) cube complex, which elucidates many structures of the groups.
Since a strongly filling surface $[S]$ separates any two points on the Gromov boundary of $G$, one can apply Bergeron-Wise boundary criterion [BW12] to construct a proper action on a CAT(0) cube complex $X$. The singleton $\{[S]\}$ satisfies the assumption of [BW12, Theorem 1.4], and thus its fundamental group $\pi_{1}(S)\rightarrow H\leq\pi_{1}(M)$ is a quasiconvex codimension-$1$ subgroup such that $G$ acts properly and cocompactly on the dual CAT(0) cube complex. The orbits of the hyperplanes correspond to the image $H$ under the natural action by $\pi_{1}(M)$. The stabilizers of the hyperplanes correspond to conjugation of $H\cong\pi_{1}(S)$.
As a consequence of Theorem 1.1, we have the following.
Corollary 3.20.
Let $G$ be the fundamental group of a closed hyperbolic $3$-manifold or a closed arithmetic hyperbolic manifold with a totally geodesic hypersurface. There exist infinitely many CAT(0) cube complexes $X$ on which $G$ acts properly, cocompactly, essentially, and hyperplane-essentially with one orbit of hyperplanes whose stabilizer is a hypersurface subgroup.
Since non commensurable strongly filling surfaces have different limit sets on $\partial G$, they give rise to cubulations that are not equivariantly isometric. This partially answers a question of Fioravanti-Hagen on [FH21, p. 879]. For any Gromov-hyperbolic group $G$ that admits a proper, cocompact action on a CAT(0) cubed complex, Fioravanti-Hagen [FH21] establish the existence of a CAT(0) cube complex $X$ on which $G$ acts properly, cocompactly, and essentially with a single orbit of hyperplanes, or acts properly, cocompactly, essentially, and hyperplane-essentially. Moreover, if we apply their theorems to the case where $G=\pi_{1}(M)$ for $M$ a closed hyperbolic $3$-manifold, it is not clear that their hyperplane-stabilizer is a surface subgroups [^2] M. Hagen informed the author that their construction applied to a filling collection of Fuchsian surfaces should give hyperplane-stabilizers that are surface subgroups. This is because their cubulation comes from taking a filling collection $A$ of finitely many surfaces constructed by [KM12] and then applying the boundary criterion [BW12]. Surfaces in $A$ may coincide on pants instead of just intersecting at closed curves. Thus it is not clear their hyperplane-stabilizers are surface subgroups. In Kahn-Wright’s construction [KW21], one uses two copies of each good pants. See also 3.16.
Fioravanti-Hagen remark on [FH21, p.880] that for certain $G$ and $X$ (including our cases), an arbitrarily small deformation of one of the original walls suffices to obtain a proper action with a single orbit of hyperplanes. It is interesting to note that using completely different techniques, i.e., homogeneous dynamics and rigidity results in [Rat91, MS95, Sha91a], we establish a similar result in our context. In particular, by Theorem 1.6, asymptotically Fuchsian surfaces are dense in $\mathcal{G}_{2}M$ and are eventually filling by Theorem 1.1. Thus in an arbitrarily small neighborhood of a geometric circle at $\partial\widetilde{M}$, there exists the limit set of a strongly filling surface.
Corollary 3.21.
Let $M$ be a closed hyperbolic $3$-manifold or a closed arithmetic hyperbolic manifold with a totally geodesic hypersurface. The collection of strongly filling hypersurfaces is ubiquitous.
3.7 Least area non-Fuchsian surfaces fail to minimize intersection numbers
Lemma 3.22.
Let $M$ be a hyperbolic $3$-manifold. Then for any least area non geodesic surface $S$, there exists a closed geodesic $\gamma$ such that the intersection number $|S\cap\gamma|$ is strictly larger than the geometric intersection number $S]\cdot[\gamma]$.
This contrasts with 3.2 and the intersection number minimizer on a hyperbolic surface.
Proof.
Take any non geodesic least area surface $S$ of genus $g$ in $M$. Since minimal surfaces in $3$-manifolds are smoothly immersed by 3.10 and non transverse self-intersections are isolated by [FHS83, Lemma 1.4], there is a point $p\in S$ not in the self-intersection locus of $S$. There exists a tangent plane at $p$ and also a totally geodesic plane $\Pi\subset M$ tangent to $S$ at $p$.
By [CM11, Theorem 7.3, Lemma 6.13] or [FHS83, Lemma 1.4], the local geometry of the minimal surface intersecting $\Pi$ can be described using classical models of non transverse intersection, which looks like $z=\textnormal{Re}(x+iy)^{n}$ intersecting $z=0$, resulting an $n$-prongs. Near $p$, there are some numbers >1>1 of parts of the minimal surface that are larger than $z=0$ (the tangent geodesic plane).
Figure 3: The least area $S$ has redundant intersections with some geodesic $\gamma$
We can draw two small open balls $B_{1},B_{2}$ in the two parts that are away from the self-intersection locus. Here we also require that $B_{1},B_{2}$ are supported in a small neighborhood $U$ of $p$, so that $U$ is away from the self-intersection locus. See Figure 3. The set of geodesics which pass through $B_{1}$ and $B_{2}$ in $M$ contains an open subset. By [KH95, Theorem 17.6.2, Corollary 6.4.20], the collection of closed geodesics is a dense subset of the unit tangent bundle of $M$. Hence, there is a closed geodesic $\gamma$ in $M$ which passes through $B_{1}$ and $B_{2}$. Then the geodesic $\gamma$ intersects the least area surface $S$ at two points which are non essential intersections, and can be removed after applying a homotopy to $\gamma$. ∎
4 From filling surfaces to the topology and geometry of the ambient manifold
Nathan M. Dunfield first suggested the existence of a filling $\pi_{1}$-injective surfaces in non hyperbolic manifolds, such as the Hantzsche-Wendt manifold. In this section, we will classify closed orientable geometric $3$-manifolds which contain a filling $\pi_{1}$-injective surface (Theorem 1.10) and prove the isoperimetric inequality between a filling surface and $M$ (Theorem 1.11).
4.1 The proof of Theorem 1.10
By 3.5, we can pass to the maximal manifolds to investigate the existence of filling surfaces. We use [AFW15, Table 1] for model geometric $3$-manifolds which are maximal among compact quotients.
Proof of Theorem 1.10.
Proof.
The spherical manifold $S^{3}$ contains a great sphere which is filling. Similarly, $\mathbb{R}P^{3}$ contains $\mathbb{R}P^{2}$ as a filling $\pi_{1}$-injective surface with nontrivial $\pi_{1}$.
Lemma 4.1.
A closed Sol $3$-manifold has no filling surfaces.
Proof.
[Mal58] proves that Sol $3$-manifolds are LERF (locally extended residually finite). If $S\hookrightarrow M$ is a $\pi_{1}$-injective filling immersed surface, by taking a finite cover $M^{\prime}$ of $M$ (which is still a torus bundle over the circle), we lift $S$ to a filling union of embedded surfaces $S^{\prime}\subset M^{\prime}$ by 3.5. In a torus bundle with Anosov monodromy, if there exists an embedded $\pi_{1}$-injective surface not homotopic to a fiber, its intersection with a fiber is a simple closed curve, which contradicts the fact the monodromy is Anosov [^3] We thank Yanqing Zou for helpful comments which simplify the proof.. Thus the only embedded essential surfaces in $M^{\prime}$ consist of fibers, which are all homotopic and whose complements are of the form $\mathbb{T}^{2}\times I$. ∎
The remaining cases are various Seifert fibered geometries. If $F$ is an immersed surface in $M$, we call it horizontal if $F$ is everywhere transverse to the fibers of $M$ and vertical if $F$ is everywhere tangent to the fibers of $M$. Any $\pi_{1}$-injective surface is homotopic to a minimal surface by [SY79, SU82]. Hass in [Has84] classifies immersed $\pi_{1}$-injective surfaces in Seifert fiber spaces using minimal surface theory.
Theorem 4.2 (Hass).
Let $M$ be a Seifert fiber space with a geometric structure. Let $S$ be a closed surface not equal to $S^{2},P^{2}$, and let $f:S\rightarrow M$ be a minimal $\pi_{1}$-injective immersion. Then $F$ is either vertical or horizontal.
A vertical surface is never filling, since the complement in $M$ contains $S^{1}$ fibers. By [Mar16, Proposition 10.4.8], if a closed Seifert fiber manifold contains a horizontal surface, its Euler class number must be 0. This rules out the Nil and the $\widetilde{\textnormal{SL}(2,\mathbb{R})}$ geometry.
Scott [Sco78] shows that the fundamental group of a Seifert fibered manifold $M$ is LERF. Thus to understand the complementary regions to an immersed $\pi_{1}$-injective surface $S$ in $M$, we can pass to a finite cover $N$ of $M$ so that the preimage of $S$ in $N$ is a finite union of embedded incompressible surfaces that are invariant under the covering group $C$. 3.5 implies that it suffices to show that no such collection of embedded incompressible surfaces fills $N$.
A closed $3$-manifold with $S^{2}\times\mathbb{R}$ geometry is finitely covered by $S^{2}\times S^{1}$. Then $S^{2}\times\theta$ is the only incompressible surface up to homotopy, and any finite union of these surfaces is non-filling.
By [Mar16, 12.3], the $6$ orientable closed Euclidean $3$-manifolds are
  1. 1.
    the $3$-torus $\mathbb{T}$ made by gluing opposite faces of a cube,
  2. 2.
    the manifold made by gluing opposite faces of a cube with a 1/2 twist on one pair,
  3. 3.
    the manifold made by gluing opposite faces of a cube with a 1/4 twist on one pair,
  4. 4.
    the manifold made by gluing opposite faces of a hexagonal prism with a $1/3$ twist on the hexagonal faces,
  5. 5.
    the manifold made by gluing opposite faces of a hexagonal prism with a $1/6$ twist on the hexagonal faces, and
  6. 6.
    the Hantzsche–Wendt manifold.
We denote these manifolds by $M_{1},\cdots,M_{6}$. Their fundamental domains are illustrated in Figure 4444This figure is taken from [Mar16, Figure 12.2], courtesy of Bruno Martelli..
Figure 4: Fundamental domains of the six closed flat orientable 33-manifolds. Unlabeled faces are glued by translations.
The $3$-torus has no filling surfaces, since every horizontal surface is a torus whose complement is $\mathbb{T}^{2}\times I$. Using ideas in 3.18 and 3.2, we prove
Lemma 4.3.
Let $M$ be a closed Euclidean $3$-manifold and $S$ a totally geodesic surface. Then the following are equivalent.
  1. 1.
    The homotopy class $[S]$ is filling.
  2. 2.
    The preimage $\widetilde{S}\subset\mathbb{R}^{3}$ contains three families of planes whose normal vectors are linearly independent.
Proof.
[…]
Readers can compare 4.3 with 3.18.
For $i=2,3,6$, $C_{i}$ denotes the covering group corresponding to the covering $M_{1}\rightarrow M_{i}$. For $i=4,5$, $C_{i}$ refers to the group corresponding to the covering from hexagonal $\mathbb{T}^{3}$ (the hexagonal fundamental domain with opposite faces glued to each other) to $M_{i}$.
The covering group $C_{2}$ of $M_{2}$ is of order $2$. The image of a torus in $\mathbb{T}$ under $C_{2}$ yields at most two families of planes in $\mathbb{R}^{3}$.
The covering groups of $M_{3},\cdots,M_{6}$ have order $\geq 3$; hence, it is not surprising that they all have filling surfaces, in view of 4.3.
The covering group $C_{3}$ of $M_{3}$ is of order $4$. A filling surface corresponds to the image of $C_{3}$ acting on the diagonal torus from the upper left edge to the lower right edge in the front face (Figure 5(a)).
For $M_{4}$, the covering group $C_{4}$ is of order $6$ and acts on the hexagonal prism (after pulling back by translation) as rotation by $60$ degrees. A filling surface corresponds to the diagonal parallelogram annulus shown in the picture Figure 5(b). The fiber directions are perpendicular to the hexagonal faces. The projection of a single annulus to the hexagonal face is a diagonal rectangle. Six of them obstruct all fibers by 4.3. The intersection of each hexagon with the six annuli consists of six arcs which cut the hexagon torus into contractible polygons. Thus the complement of the surface in $M_{4}$ is contractible.
(a) The annulus under $90^{\circ}$ rotation generates a filling surface in $M_{3}$
(b) The annulus under $60^{\circ}$ rotation generates a filling surface in $M_{4}$
(c) The annulus under $120^{\circ}$ rotation generates a filling surface in $M_{5}$
Figure 5: Filling surfaces in flat manifolds
Similarly, we construct a filling surface in $M_{5}$ (Figure 5(c)).
The last one is the Hantzsche-Wendt manifold $M_{6}$, which has a $4$-fold cover the $3$-torus $M_{1}$. From the isometries at [CM22, page 3], the covering group $C_{6}$ acting on some torus $S$ (chosen later) results in four copies of $S$ each representing a different homology class in $H_{2}(\mathbb{T}^{3};\mathbb{Z})$. Denote the projection by $\pi_{6}:M_{1}\rightarrow M_{6}$. Take $S$ to be a torus with normal vector $(1,1,1)$, then the other the other tori have normals (1,1,1),(1,1,1)(-1,-1,1),(-1,1,-1), and (1,1,1)(1,-1,-1), resulting in contractible complementary regions by 4.3. Thus $\bar{S}\coloneqq\pi_{6}(S)$ is a filling surface in $M_{6}$.
A closed $3$-manifold with $\mathbb{H}^{2}\times\mathbb{R}$ geometry has a finite cover $S_{g}\times S^{1}$. We adapt the construction of a filling surface in $M_{4}$. For example, consider instead of a hexagon prism an octagon prism $O\times I$ with proper gluings to form $S_{2}\times S^{1}$. A quotient manifold $M$ is the prism with $O\times 0$ and $O\times 1$ identified by a translation with a $2\pi/8$-twist ($S_{2}\times S^{1}\rightarrow M$ is an $8$-fold cover). The annulus formed by gluing two parallelograms in Figure 6 generates a filling surface $S$ in $M$ under the covering group. Now we show $[S]$ is filling. The annulus generates $16$ families of planes in $\mathbb{H}^{2}\times\mathbb{R}$. Let $\tilde{\gamma}$ be a vertical fiber perpendicular to the octagon face $O\times 1$. Then $\tilde{\gamma}$ meets one of the planes exactly once, implying $\pi_{1}(M-S)$ does not contain $\gamma]$. Similarly, if $\tilde{\gamma}$ is any other geodesic, then we can find a plane $\widetilde{P}$ from the $16$ planes so that $\tilde{\gamma}$ and $\widetilde{P}$ intersect exactly once. This implies that $[S]$ is filling.
Figure 6: A filling surface in an $\mathbb{H}^{2}\times\mathbb{R}$ manifold
4.2 The measures and ranks of a filling surface and its ambient manifold
The proof of Theorem 1.11
Proof.
Suppose $[S]$ is filling. By 2.7, a least area representative $S$ cuts $M$ into balls. The boundary components of MSM-S together form two copies of $S$. By [Rat19, (3.4.6)], the hyperbolic metric in spherical coordinates is
[…]
Then the volume of a hyperbolic $3$-ball is
(4.1) $\textnormal{Vol}(B(r))=\textnormal{Area}(S^{2})\int_{0}^{r}\sinh^{2}t\,\,\mathrm{d}t=\frac{\pi}{2}(e^{2r}-e^{-2r}-4r)$
while the hyperbolic area
(4.2) $\textnormal{Area}(\partial B(r))=\textnormal{Area}(S^{2})\sinh^{2}r=\pi(e^{2r}+e^{-2r}-2).$
It is elementary to prove $2e^{-2r}+4r\geq 2$ and hence $\textnormal{Area}(\partial B(r))\geq 2\textnormal{Vol}(B(r)).$ Now consider a ball $B(r)$ such that $\textnormal{Vol}(B(r))=\textnormal{Vol}(M)$. By classical isoperimetric inequality for topological balls in hyperbolic $3$-space (e.g., [Coo99, p.2]), we have
$\textnormal{Area}(M-S)=2\textnormal{Area}(S)\geq\textnormal{Area}(\partial B(r))\geq 2\textnormal{Vol}(B(r))=2\textnormal{Vol}(M).$
Equality is only achieved at $r=0$ and hence it is a strict inequality. Let $V(r)=\textnormal{Vol}(B(r))$. [BGLS10, p. 2204] gives an explicit formula relating the rank of a lattice to its covolume, which for $PSL_{2}(\mathbb{C})$ is
$d(\Gamma)\leq\frac{V(1.25\epsilon)}{V(0.25\epsilon)^{2}}\textnormal{Vol}(M)<1.698\times 10^{9}\textnormal{Vol}(M),$
where $\epsilon=0.104/10$ (we have used Meyerhoff’s estimate [Mey87] of the Margulis constant) and we have used (4.1). ∎
Example 4.4.
There exists a sequence of $3$-manifolds $M_{j}$ admitting a filling torus such that $\textnormal{rank}(M_{j})\rightarrow\infty$.
Proof.
The construction in Figure 6 is an immersed annulus which under the covering transformations glues to a filling torus in the manifold with an eight-fold cover $S_{2}\times S^{1}$. The same construction applies to the manifold whose $4g$-fold cover is $S_{g}\times S^{1}$ and produces a filling torus. ∎